# GoRINNs: Godunov-Riemann Informed Neural Networks for Learning Hyperbolic Conservation Laws

### v2

@@ Abstract
We present GoRINNs: numerical analysis-informed neural networks for the solution of inverse problems of non-linear systems of conservation laws. GoRINNs are based on high-resolution Godunov schemes for the solution of the Riemann problem in hyperbolic Partial Differential Equations (PDEs). In contrast to other existing machine learning methods that learn the numerical fluxes of conservative Finite Volume methods, GoRINNs learn the physical flux function per se. Due to their structure, GoRINNs provide interpretable, conservative schemes, that learn the solution operator on the basis of approximate Riemann solvers that satisfy the Rankine-Hugoniot condition. The performance of GoRINNs is assessed via four benchmark problems, namely the Burgers’, the Shallow Water, the Lighthill-Whitham-Richards and the Payne-Whitham traffic flow models. The solution profiles of these PDEs exhibit shock waves, rarefactions and/or contact discontinuities at finite times. We demonstrate that GoRINNs provide a very high accuracy both in the smooth and discontinuous regions.
@@ 1 Introduction
Conservation laws (of mass, momentum, and energy) are fundamental physical laws present in many problems including fluid mechanics [4], traffic flow and crowd dynamics [74, 5], electromagnetism [53], among others, in which a physical quantity is conserved in time over an isolated spatial domain [41, 18]. The conserved quantities are physically determined or can arise in complex systems at the macroscopic level from the microscopic interactions (particles, molecules, vehicles, individuals, etc); see e.g., [11]. Conservation laws are often modelled with hyperbolic, (systems of) Partial Differential Equations (PDEs) [41], which can exhibit the challenging feature of emerging discontinuities, such as shock waves, at finite times even when the initial conditions are smooth [18].
@@
The solution of the so-called forward problem, has been the objective of many Scientific Machine Learning (SciML) algorithms. One broad category consists of physics-informed deep neural networks (PINNs) [62, 55, 81, 52] which however, when employed for hyperbolic PDEs, have a rather poor performance/low accuracy near the shock waves [50, 57]. To tackle this challenge, several PINN-based schemes have been proposed for respecting the conservation law [57, 82, 19, 63]. Finally, other methods include hybrid methods, which combine the classical numerical methods (to guarantee conservation) with neural networks (NNs) [77, 37, 49, 68, 76].
@@
For the solution of the inverse problem (that is learning functional terms/closures/boundary conditions of unknown PDEs, as well as values of the parameters from available data), SciML methods typically include estimation of the time and spatial derivatives from time-series data, which are then provided in sparse regression models [69, 67, 34], DNNs [40, 25] and Gaussian Process Regressors [39, 61]. However, such schemes face significant challenges when employed to find closures of hyperbolic PDEs, since they do not preserve explicitly the conservation laws. In fact near shocks, the finite difference schemes for the estimation of the derivatives cannot accurately learn the true differential operator, and NN-based differentiation schemes may be required [3]. In addition, even when the derivative estimation is accurate, the above methods may introduce artificial differential terms in the related closures (see e.g. inviscid Burgers’ equation in [69]). An alternative approach is the PINNs and their conservative flavors [55, 81, 52, 50, 32, 57, 1]. Neural Operators (NOs) such as DeepONets [46], Fourier Neural Operators [43, 38, 72], and RandONets [24], have been also introduced for learning the evolution operator, i.e. the right hand side of PDEs and/or the solution operator. However, NOs are also prone to discontinuities, and only few works are dedicated to hyperbolic PDEs [79, 38]. Recently, RiemannONets [60], a special variation of DeepONets [46], were introduced to approximate with high accuracy discontinuous solutions of the Riemann problem for hyperbolic equations.
@@
While such schemes are highly accurate for parameter estimation tasks, they do not explicitly provide conservative schemes, nor have not so far been employed, to the best of our knowledge, for learning the functional closures in PDEs of conservation laws. For learning the flux functionals, hybrid methods have been recently developed, combining SciML algorithms with numerical methods [54, 12, 36]. The above hybrid methods are especially crafted for respecting conservation laws. However, they, learn the numerical approximation of the flux rather than the physical one. This may introduce significant inaccuracy for coarsely discretized data. For learning the physical, parameterized, flux, in [42], a special network architecture, resulting to rationals of polynomial functions, was coupled with the Rusanov (local Lax-Friedrichs) numerical FV scheme. Leaving aside the accuracy of the selected FV scheme (which is usually of first-order), all the aforementioned hybrid methods have been developed for non-linear scalar hyperbolic PDEs, in which the discontinuities (shocks) usually travel along one characteristic. Clearly, this is not the case for systems of hyperbolic PDEs.
@@
In this work, we present a hybrid approach for the solution of the inverse problem for both scalar and systems of hyperbolic PDEs in 1-dim. spatial domains. In particular, we couple NNs with high-resolution Godunov-type FV schemes for the solution of the Riemann problem. We name this scheme as Godunov-Riemann Informed Neural Networks (GoRINNs).
@@
In contrast to existing hybrid methods that discover numerical fluxes [54, 12, 36], here we use shallow NNs to approximate directly the physical flux functional terms, which are missing from the flux function of the unknown system of PDEs. To take into account partial information that may be available, GoRINNs can be also employed in the presence of known fluxes or source terms in the system of PDEs. For obtaining the numerical solution of the unknown system, we perform operator splitting and solve the homogeneous part of the PDEs with high-resolution Godunov-type methods with flux-limiters and the inhomogeneous part (if any) with stiff integration schemes. These numerical schemes provide high-order accuracy both far and near the emerging discontinuities. However, since we deal with non-linear systems of hyperbolic PDEs, the related Riemann problem is solved by approximate Riemann solvers [41], which need to satisfy the Rankine-Hugoniot (RH) condition in order for the numerical scheme to be conservative [73, 41]. While NNs have been used to obtain accurate and RH condition-satisfying approximations for the solution of the Riemann problem via regression [49, 68, 76], here we solve the inverse problem with unknown flux terms; hence the target solutions are simply not available. To overcome this obstacle, we construct Riemann solvers that are dependent on the unknown flux function and constrain them to satisfy the RH condition; the latter is implemented as a soft constraint.
@@
The performance of GoRINNs is assessed via four benchmark problems of non-linear conservation laws arising in fluid dynamics and traffic flow. First, we consider the scalar conservation laws of the inviscid Burgers’ equation [9] and the Lighthill-Whitham-Richards (LWR) equation [44, 64], for which we learn the flux function and the velocity closure, respectively. Next, we study two systems of two conservation laws, the Shallow Water (SW) and the Payne-Whitham (PW) equations [59, 80], for both of which we learn the pressure closures in the flux function. We highlight here, that the Burgers’, LWR and SW equations are homogeneous hyperbolic PDEs that exhibit analytic solutions for the Riemann problem, while for the PW equations not only the Riemann problem cannot be solved analytically, but also source terms are included. For all the benchmark problems, the approximate Riemann solvers are derived in appendix C. We underline here that for all problems under study, GoRINNs do not require extensive training data sets for achieving high approximation accuracy. This is because, a low number of complete (in space), time observations results in a high number of residuals to be minimized for the optimization problem, even for coarse spatial discretizations.
The reminder of the paper is organized as follows. In section 2, we concisely describe the problem definition. In section 3, we briefly present the high-resolution Godunov-type FV numerical scheme and introduce in detail GoRINNs for the solution of the inverse problem. In section 4, we present the numerical results for the four benchmark problems considered (which are presented in Appendix A along with details about the forward problem solution required for the data acquisition) and assess the convergence and accuracy of GoRINNs. Finally, in section 5, we conclude, discussing limitations and perspectives of the present work.
@@ 5 Conclusions
Hyperbolic partial differential equations (PDEs) that model conservation laws play a central role in explaining complex phenomena across diverse disciplines such as fluid mechanics, aerospace engineering, electromagnetics, traffic flow, and crowd dynamics [5]. The numerical solution of the forward problem for such PDEs is challenging due to the presence of discontinuities, such as shock waves and rarefactions, which can cause classical numerical methods to fail or yield inaccurate solutions. The inverse problem — i.e., that of learning the underlying closures/potentials or values of parameters in these PDEs from observed data — poses even greater difficulties as the problem is often ill-posed, thus relatively small approximation errors can lead to instabilities and violation of the conservation laws.
@@
In the last few years, SciML deep-learning methods, such as PINNs [55, 81, 50, 57, 32] and neural operators [46, 79, 38, 72], have shown great potential in dealing with the challenging inverse problem of parameter inference in conservation laws. One critical issue is the lack of interpretability [60, 21, 24]. By construction most of the until now proposed schemes, set the conservation law as a soft constraint in the loss function, thus failing to respect accurately the conservation property. By explicitly preserving the conservation laws, hybrid SciML methods infer fluxes of conservative FV schemes [54, 12, 36], which however provide representation of the numerical-approximated quantities rather than the physical ones per se. These methods may introduce significant inaccuracies for coarsely discretized data and are developed for scalar conservation laws.
@@
GoRINNs offer interpretable shallow feedforward NNs for tackling the above problems. Unlike other hybrid SciML methods that approximate the numerical fluxes, GoRINNs learn the solution of the Godunov scheme by directly approximating the physical flux functions per se. This approach enhances interpretability, thus respecting explicitly the underlying conservation laws as approximate Riemann solvers that satisfy the Rankine-Hugoniot condition. Moreover, GoRINNs provide a versatile framework for identifying flux functionals directly from data. This approach accommodates systems of conservation laws and allows for the inclusion of source terms—whether known or unknown—in the associated hyperbolic PDEs. As we show, GoRINNs are able to approximate accurately the closures of four benchmark problems, which exhibit traveling shocks that emerge at finite times. Based on the solution of the inverse problem, which actually approximates the physical flux function required for the solution operator of the Godunov-type scheme, we solved the forward problem for long time periods, showing that GoRINNs result to a high numerical approximation accuracy, preserving numerical stability and the conservation laws.
@@
The training and performance of GoRINNs, like other SciML methods, benefit from partial knowledge of the physics underlying the missing terms (see section 4.4, where GoRINNs converge with the “correct” predictor selection.) If no physical insight is available, then techniques such as symbolic regression [75, 6], manifold learning (e.g., parsimonious diffusion maps) [15, 22, 20, 25, 56], and SINDy [8, 10] can be exploited to learn a parsimonious set of variables that can parametrize the unknown closures. Furthermore, GoRINNs can be extended to analyze real-world data derived from microscopic simulations/observations, facilitating the discovery of macroscopic hyperbolic PDEs for the emergent dynamics. This application will further require parameter inference, a task that can be addressed by including physical parameters as additional GoRINNs predictors. Finally, using GoRINNs to discover macroscopic closures connects directly to control theory, which can be exploited for enabling distributed control in large-scale microscopic models (see e.g., [47, 48]), given GoRINNs-derived information at the macroscopic level.

### v3

@@ Abstract
We present GoRINNs: numerical analysis-informed (shallow) neural networks for the solution of inverse problems of non-linear systems of conservation laws. GoRINNs is a hybrid/blended machine learning scheme based on high-resolution Godunov schemes for the solution of the Riemann problem in hyperbolic Partial Differential Equations (PDEs). In contrast to other existing machine learning methods that learn the numerical fluxes or just parameters of conservative Finite Volume methods, relying on deep neural networks (that may lead to poor approximations due to the computational complexity involved in their training), GoRINNs learn the closures of the conservation laws per se based on “intelligently“ numerical-assisted shallow neural networks. Due to their structure, IN particular, GoRINNs provide explainable, conservative schemes, that solve the inverse problem for hyperbolic PDEs, on the basis of approximate Riemann solvers that satisfy the Rankine-Hugoniot condition. The performance of GoRINNs is assessed via four benchmark problems, namely the Burger’s, the Shallow Water, the Lighthill-Whitham-Richards and the Payne-Whitham traffic flow models. The solution profiles of these PDEs exhibit shock waves, rarefactions and/or contact discontinuities at finite times. We demonstrate that GoRINNs provide a very high accuracy both in the smooth and discontinuous regions.
@@ 1 Introduction
Conservation laws (of mass, momentum, and energy) are fundamental physical laws present in many problems including fluid mechanics including multiphase problems [4, 26], traffic flow and crowd dynamics [78, 5], electromagnetism [55], collisionless shocks for astrophysical phenomena[79] among others, in which a physical quantity is conserved in time over an isolated spatial domain [43, 18]. The conserved quantities are physically determined or can arise in complex systems at the macroscopic level from the microscopic interactions (particles, molecules, vehicles, individuals, etc); see e.g., [11]. Conservation laws are often modelled with hyperbolic, (systems of) Partial Differential Equations (PDEs) [43], which can exhibit the challenging feature of emerging discontinuities, such as shock waves, at finite times even when the initial conditions are smooth [18].
@@
The solution of the so-called forward problem, has been the objective of many Scientific Machine Learning (SciML) algorithms. One broad category consists of physics-informed deep neural networks (PINNs) [66, 59, 86, 54] which however, when employed for hyperbolic PDEs, have a rather poor performance/low accuracy near the shock waves [52, 61]. To tackle this challenge, several PINN-based schemes have been proposed for respecting the conservation law [61, 87, 19, 67]. Finally, other methods include hybrid/ blended methods, which combine the classical numerical methods (to guarantee conservation) with neural networks (NNs) [82, 38, 39, 51, 58, 72, 81].
@@
For the solution of the inverse problem (that is learning functional terms/closures/boundary conditions of unknown PDEs, as well as values of the parameters from available data), SciML methods typically include estimation of the time and spatial derivatives from time-series data, which are then provided in sparse regression models [73, 71, 35], DNNs [42, 25] and Gaussian Process Regressors [41, 65]. However, such schemes, relying mostly on DNNs, face significant challenges when employed to find closures of hyperbolic PDEs, since they do not preserve explicitly the conservation laws, thus providing rather moderate accuracy due to the computational complexity involved in the training of DNNs. In fact near shocks, the finite difference schemes for the estimation of the derivatives cannot accurately learn the true differential operator, and NN-based differentiation schemes may be required [3]. In addition, even when the derivative estimation is accurate, the above methods may introduce artificial differential terms in the related closures (see e.g. inviscid Burgers’ equation in [73]). Approaches, relying to physics-informed neural networks (PINNs) and their conservative flavors [59, 86, 54, 52, 33, 61, 1]. Neural Operators (NOs) such as DeepONets [48], Fourier Neural Operators [45, 40, 76], and RandONets [24], have been also introduced for learning the evolution operator, i.e. the right-hand side of PDEs and/or the solution operator. However, NOs are also prone to discontinuities, and only few works are dedicated to hyperbolic PDEs [84, 40] whle the problem of computational complexity remains. Recently, RiemannONets [64], a special variation of DeepONets [48], were introduced to approximate with high accuracy discontinuous solutions of the Riemann problem for hyperbolic equations. For a review on the solution for the inverse problem for PDEs see also in [58].
@@
While such schemes have been used for parameter estimation tasks, they do not explicitly provide conservative schemes, nor have not so far been employed, to the best of our knowledge, for learning the functional closures in PDEs of conservation laws. For learning the flux functionals/closures, hybrid/blended methods have been also developed, combining SciML algorithms with numerical methods [58, 57, 12, 37]. The above hybrid methods are especially crafted for respecting conservation laws. However, they, learn the numerical approximation of the fluxes/closures rather than the physical closures per se. This may introduce significant inaccuracy for coarsely discretized data. For learning the physical, parameterized, flux closure, in [44], a special network architecture, resulting to rationals of polynomial functions, was coupled with the Rusanov (local Lax-Friedrichs) numerical FV scheme. Leaving aside the numerical approximation accuracy of the selected FV scheme (which is usually of first-order), all the aforementioned hybrid methods have been developed for non-linear scalar hyperbolic PDEs, in which the discontinuities (shocks) usually travel along one characteristic. Clearly, this is not the case for systems of hyperbolic PDEs.
@@
In this work, we present a hybrid/blended approach for the solution of the inverse problem for both scalar and systems of hyperbolic PDEs in 1-dim. spatial domains. In particular, we couple shallow NNs with high-resolution Godunov-type FV schemes for the solution of the Riemann problem. We name this scheme as Godunov-Riemann Informed Neural Networks (GoRINNs).
@@
In contrast to existing hybrid methods that discover parameters or numerical fluxes [57, 12, 37], based on DNNs, here we use “intelligently“ designed numerical analysis informed shallow NNs to approximate directly the physical flux functional terms, which are missing from the flux function/closure of the unknown system of PDEs. To take into account partial information that may be available, GoRINNs can be also employed in the presence of known physics such as known fluxes or source terms in the system of PDEs. For obtaining the numerical solution of the unknown system, we perform operator splitting and solve the homogeneous part of the PDEs with high-resolution Godunov-type methods with flux-limiters and the inhomogeneous part (if any) with stiff integration schemes. These numerical schemes provide high-order accuracy both far and near the emerging discontinuities. However, since we deal with non-linear systems of hyperbolic PDEs, the related Riemann problem is solved by approximate Riemann solvers [43], which need to satisfy the Rankine-Hugoniot (RH) condition in order for the numerical scheme to be conservative [77, 43]. While NNs have been used to obtain RH condition-satisfying approximations for the solution of the Riemann problem via regression [51, 72, 81], here we solve the inverse problem with unknown flux terms/closures; hence the target solutions are simply not available. To deal with this problem, we construct Riemann solvers that are dependent on the unknown flux function/closure and constrain them to satisfy the RH condition; the latter is implemented as a soft constraint.
@@
The performance of GoRINNs is assessed via four benchmark problems of non-linear conservation laws arising in fluid dynamics and traffic flow. First, we consider the scalar conservation laws of the inviscid Burgers’ equation [9] and the Lighthill-Whitham-Richards (LWR) equation [46, 68], for which we learn the flux function and the velocity closure, respectively. Next, we study two systems of two conservation laws, the Shallow Water (SW) and the Payne-Whitham (PW) equations [63, 85], for both of which we learn the pressure closures in the flux function. We highlight here, that the Burgers’, LWR and SW equations are homogeneous hyperbolic PDEs that exhibit analytic solutions for the Riemann problem, while for the PW equations not only the Riemann problem cannot be solved analytically, but also source terms are included. For all the benchmark problems, the approximate Riemann solvers are derived in C. We underline here that for all problems under study, GoRINNs due to their structure do not require extensive training data sets for achieving high approximation accuracy. This is because, a low number of complete (in space), time observations results in a high number of residuals to be minimized for the optimization problem, even for coarse spatial discretizations.
The reminder of the paper is organized as follows. In section 2, we concisely describe the problem definition. In section 3, we briefly present the high-resolution Godunov-type FV numerical scheme and introduce in detail GoRINNs for the solution of the inverse problem. In section 4, we present the numerical results for the four benchmark problems considered (which are presented in A along with details about the forward problem solution required for the data acquisition) and assess the convergence and accuracy of GoRINNs. Finally, in section 5, we conclude, discussing limitations and perspectives of the present work.
@@ 5 Conclusions
Hyperbolic partial differential equations (PDEs) that model conservation laws play a central role in explaining complex phenomena across diverse disciplines such as fluid mechanics, aerospace and aerodynamics engineering, acoustics, astrophysics, electromagnetics, traffic flow, and crowd dynamics, to name just a few. The numerical solution of the forward problem for such PDEs is challenging due to the presence of discontinuities, such as shock waves and rarefactions, which can cause classical numerical methods to fail or yield inaccurate solutions. The inverse problem —especially that of learning the physical closures/potentials — poses even greater difficulties because of the inherent nonlinearities, ill-posedness, and emerging discontinuities. Thus, relatively small approximation errors can lead to instabilities and violation of the conservation laws.
@@
In the last few years, SciML based on DNNs, such as PINNs [59, 86, 54, 52, 61, 33] and neural operators [48, 45, 84, 40, 76], have shown great potential in dealing with the challenging inverse problem of parameter inference in conservation laws. One critical issue is the lack of explainability [64, 21, 24] and the computational complexity pertaining to their training, thus results to rather moderate approximation accuracy (see also the discussion and some demonstrations in [24]). By construction, most of the until now proposed schemes, set the conservation law as a soft constraint in the loss function, thus failing to respect accurately the conservation property. By explicitly preserving the conservation laws, hybrid SciML methods infer fluxes of conservative FV schemes [57, 12, 37], which however provide representation of the numerical-approximated quantities rather than the physical ones per se. These methods may introduce significant inaccuracies for coarsely discretized data and are developed for scalar conservation laws.
@@
GoRINNs offer explainable “intelligently” designed shallow feedforward NNs for tackling the above problems. Unlike other hybrid SciML methods that approximate the numerical fluxes using DNNs, GoRINNs solve the inverse problem enhancing explainability, and accuracy, thus respecting explicitly the underlying conservation law as approximate Riemann solvers that satisfy the Rankine-Hugoniot condition. This approach accommodates systems of conservation laws and allows for the inclusion of source terms—whether known or unknown—in the associated hyperbolic PDEs. As we show, GoRINNs are able to approximate with high accuracy the closures of four benchmark problems, which exhibit traveling shocks that emerge at finite times. We show that GoRINNs result to a high numerical approximation accuracy, preserving numerical stability and the conservation laws.
@@
The training and performance of GoRINNs, can also incorporate partial knowledge of the physics underlying the missing terms (see section 4.4, where GoRINNs converge with the “correct” predictor selection). If no physical insight is available, then techniques such as symbolic regression [80, 6], manifold learning (e.g., parsimonious diffusion maps) [15, 22, 20, 25, 60], and SINDy [8, 10] can be exploited to learn a parsimonious set of variables that can parametrize the unknown closures. Furthermore, GoRINNs can be extended to analyze real-world data derived from microscopic simulations/observations, facilitating the discovery of macroscopic hyperbolic PDEs for the emergent dynamics. This application will further require parameter inference, a task that can be addressed by including physical parameters as additional GoRINNs predictors. Finally, using GoRINNs to discover macroscopic closures connects directly to control theory, which can be exploited for enabling distributed control in large-scale microscopic models (see e.g., [49, 50]), given GoRINNs-derived information at the macroscopic level.
Importantly, what we propose and highlight here is the need of developing new SciML resource-bounded algorithms that can provide high accuracy at a low computational cost. That is develop new “intelligently” designed schemes which are not only physics-informed but also numerical analysis-informed in order to deal with the time and space complexity for training DNNs. Towards this aim, the recently introduced blended inverse-PDE networks (BiPDE-Nets) [58], first discover unknown terms of PDEs and in a latter stage carry out numerical operations similar to those in traditional solvers. In [56], the authors show that “accurate” optimization of shallow NNs can yield superior results compared to training DNNs for PDEs. Recently, RandONets [24] have been introduced to learn linear and nonlinear operators based on random projections, shallow NNs, and iterative linear algebra methods for (large-scale) ill-posed problems. There, it has been shown that RandONets outperform DNNs by several orders of magnitude both in terms of computational cost and accuracy, reaching also machine-precision accuracy for linear operators. Our hybrid/blended scheme presented here, contributes exactly towards this important direction of research in SciML, focusing on the solution of the inverse problem for hyperbolic conservation laws with important applications across many fields.